OpenStudy (anonymous):

how do i find a simpler function that agrees to this function? lim as x approaches 0 for e^2x-1/e^x-1

OpenStudy (anonymous):

factor and cancel

OpenStudy (anonymous):

\[\frac{x^2-1}{x-1}=x+1\]

OpenStudy (anonymous):

how exactly did you factor?.

OpenStudy (aum):

\(e^{2x} - 1 = (e^x)^2 - 1^2 = (e^x+1)(e^x-1)\)

OpenStudy (anonymous):

oh ! okay, then x-1 cancels out and leaves you with x+1 ?

OpenStudy (aum):

\(e^x-1\) cancels out leaving you with \(e^x+1\)

OpenStudy (anonymous):

i confused you sorry \[\frac{e^{2x}-1}{e^x-1}=\frac{(e^x-1)(e^x+1)}{e^x-1}=e^x+1\]

OpenStudy (anonymous):

oh so then the limit would =1 right?

OpenStudy (anonymous):

\[e^0+1=?\]

OpenStudy (anonymous):

wait sorry 2

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you!